Using simple interest on a loan at 8.4% for one year making monthly payments. Find the APR on a $5000 loan. Write as a decimal rounded to two values.
Answer #1

Principle amount = 5000

Simple interest rate = 8.4%

Yearly interest amount = 5000*8.4/100 = 420

Annual Percentage Rate (APR) = annual rate of interest charged

= percentage of actual yearly cost of funds over the principle amount

= (Fees + annual interest)*100/principle amount = 420*100/5000 = 8.4%